Help with these problems due tomorrow night. Thank you
1. Solve by factoring:
x2 – 6x = -8


2. Solve using the square root property:
2x2 – 3 = 15


3. Solve using the square root property:
(x + 7)2 = 81


4. Solve by completing the square:
x2 + 8x + 2 = 0



Answer by Nate(3500)

1. Solve by factoring:
x2 – 6x = -8
x^2 - 6x + 8 = 0
x^2 - 4x - 2x + 8 = 0
(x^2 - 4x) + (-2x + 8) = 0
x(x - 4) - 2(x - 4) = 0
(x - 2)(x - 4) = 0
x = 2 and x = 4
2. Solve using the square root property:
2x2 – 3 = 15
2x^2 = 18
x^2 = 9
x = +- 3
3. Solve using the square root property:
(x + 7)2 = 81
x + 7 = +- 9
x = -7 +- 9
x = -16 or x = 2
4. Solve by completing the square:
x2 + 8x + 2 = 0
x^2 + 8x = -2
(x + 4)^2 = 14
x + 4 = +- sqrt(14)
x = -4 +- sqrt(14)